MONT 104N – Modeling the Environment Third “Debates” Day November 28, 2012

Presenters

Chung Nguyen, Caroline O’Day, Mike Ortlieb, Anabella Perez, Chris Simpson

The Human Population of Planet Earth

Many of the modeling techniques we have discussed have been applied to questions about human populations as well as animal and plant populations. Understanding what will happen to the human population of the Earth over the next several centuries is a critical component of understanding and controlling our impact on the planet and its ecosystems, climate, and so forth.

There is an interesting collection of positions staked out in the New York Times Room For Debate blog posted on May 4, 2011 (this appeared before the best estimates for the total human population reached 7 billion in October, 2011 – a milestone, to say the least!):

              http://www.nytimes.com/roomfordebate/2011/05/04
                 /can-the-planet-support-10-billion-people

(this should be all entered as one web address; it’s broken over two lines to fit on the page here) including several touching directly on issues we are considering in this course.

Assignment

For this “debate,” I think it will again be a good plan for each of you to take the position of one of the six contributors to the New York Times blog, look into who your people are, what their backgrounds are, why they might hold the opinions they do, and then research some of the details behind what they are saying.

Everyone can use the computer and projector in our classroom to display any graphical or numerical information they would like to bring in.

The natural way to do this will be to have each of you take one position and plan for each person to speak for about 4 or 5 minutes, with time for a follow-up discussion after everyone has spoken. Here is a rough idea about how to divide things up

1. Chung – Jamais Cascio, How 10 Billion Can Survive

2. Caroline – Jason Clay, More Efficient Food Production

3. Mike – Joel Cohen, We Can Change the Future

4. Anabella – Warren Sanderson, Educate the Masses

5. Chris – Brad Allenby, Technology’s Impact on Population (Note that I am leaving out the post on the specific problems of the continent of Africa. If you want to use some of the ideas presented there for examples, though, feel free.)

I would like to meet with the whole team before class on Wednesday, December 5 to see where you are and see if you have any questions about how to make your cases. I’m free basically during office hours 10am - 12 noon on Monday, December 3 and all morning on Tuesday, December 4.